# Exploring Scene Affinity for Semi-Supervised LiDAR Semantic Segmentation

### v2

@@ Abstract
This paper explores scene affinity (AIScene), namely intra-scene consistency and inter-scene correlation, for semi-supervised LiDAR semantic segmentation in driving scenes. Adopting teacher-student training, AIScene employs a teacher network to generate pseudo-labeled scenes from unlabeled data, which then supervise the student network’s learning. Unlike most methods that include all points in pseudo-labeled scenes for forward propagation but only pseudo-labeled points for backpropagation, AIScene removes points without pseudo-labels, ensuring consistency in both forward and backward propagation within the scene. This simple point erasure strategy effectively prevents unsupervised, semantically ambiguous points (excluded in backpropagation) from affecting the learning of pseudo-labeled points. Moreover, AIScene incorporates patch-based data augmentation, mixing multiple scenes at both scene and instance levels. Compared to existing augmentation techniques that typically perform scene-level mixing between two scenes, our method enhances the semantic diversity of labeled (or pseudo-labeled) scenes, thereby improving the semi-supervised performance of segmentation models. Experiments show that AIScene outperforms previous methods on two popular benchmarks across four settings, achieving notable improvements of 1.9% and 5.3% in the most challenging 1% labeled data.
@@ 1 Introduction
AIScene has been extensively evaluated through experiments and demonstrated superiority on two popular benchmarks for semi-supervised LiDAR semantic segmentation. It exceeds state-of-the-art methods by 1.3% on the SemanticKITTI [2] and 2.4% on the nuScenes [10] in terms of average performance across four settings (1%, 10%, 20%, and 50% labeled data). Notably, in the most challenging setting, with only 1% labeled data, AIScene achieves remarkable improvements of 1.9% and 5.3% on the two datasets. Additionally, our ablation studies confirm AIScene’s potential to reduce the requirements for labeled data across different LiDAR semantic segmentation networks (see Tab. 1 and  Tab. 2), with each component of AIScene able to function independently to facilitate semi-supervised LiDAR semantic segmentation (see Fig. 6).
@@ 2 Related Work
3D Point Cloud Data Augmentation: Point cloud data augmentation has been explored across various 3D vision tasks [54, 33, 42, 9]. Inspired by the success of Mixup [49] in image classification, several methods have investigated techniques for mixing selected point cloud samples. To preserve structural information, RSMix [20] introduces a neighboring function, and PointCutMix [50] establishes correspondences for sample fusion. Notably, these methods focus on local object-level augmentation. In response, several works have explored mixing strategies for point cloud scene augmentation. PloarMix [45] swaps points within the fan-shaped areas between two scenes, while CoSMix [31] aims to mitigate domain shift by mixing semantic point groups. Unlike these methods, our method is capable of modeling inter-scene information across multiple scenes to create more diversified data, while also extracting instance-level semantic information for individual classes to further enrich scenes.
@@ 3.3 Patch-based Data Augmentation
Following the same steps used for labeled scenes, we further create a pseudo patch pool $\mathcal{B}_{\text{Pat}}^{u}$ and a pseudo instance pool $\mathcal{B}_{\text{Ins}}^{u}$ from the erased scene $\hat{x}_{i}^{u}$. Unlike $\mathcal{B}_{\text{Pat}}^{l}$ and $\mathcal{B}_{\text{Ins}}^{l}$ covering the whole training set $D_{l}$, the pseudo pools $\mathcal{B}_{\text{Pat}}^{u}$ and $\mathcal{B}_{\text{Ins}}^{u}$ only contain scene and instance information from the current batch and are reset in each iteration. This prevents low-quality pseudo labels from persisting in pools and misleading model training. Additionally, if the number of points within a patch or instance falls below a certain threshold $\tau_{\text{min}}$, it will not be saved into the pool, ensuring effective semantic diversity.
@@
Ours (AIScene) - 61.2 66.3 67.4 67.9 65.7 63.4 73.7 75.0 76.3 72.1
Improvement $\uparrow$ - +1.9 +1.2 +1.1 +0.9 +1.3 +5.3 +3.5 +1.0 -0.2 +2.4
Table 1: Experimental results on the SemanticKITTI and nuScenes dataset compared with different SSL methods. The results are reported with mIoU scores (%). The blod and underline scores respectively denote the best and second results.
@@ 4.2 Implementation Details
We conduct experiments with 4$\times$ NVIDIA V100 GPUs with 32GB RAM, based on MMDetection3D [8]. Inspired by the GT-sampling strategy [47], we set $\tau_{min}$ to 5 in our experiments. For the threshold $\tau_{s}$ for pseudo-label generation and point erasure, we use the same parameter setting as in LaserMix. We use the MinkowskiNet [7] and Cylinder3D [55] as the backbone network due to its widely adopted in previous methods [15, 27, 41]. To facilitate patch-based data augmentation, we restrict our focus to point clouds within the X and Y dimensions in the range of $[-50,50]m$ for both the SemanticKITTI and nuScenes datasets. Similar to mainstream work, we also adopt the consistency loss proposed in MeanTeacher [38] to align the predictions of the teacher and student networks. Additionally, the weights of the consistency loss are assigned as $1\times 10^{3}$ and $2\times 10^{3}$ for the SemanticKITTI and nuScenes datasets, respectively. The weights for $\lambda_{u}$ and $\lambda_{l}$ are set as 1 for both datasets. Our method excludes all forms of model ensemble and test-time augmentation during the evaluation.
@@ 4.3 Main Results
Tab. 1 summarizes the quantitative results on SemanticKITTI and nuScenes datasets. Our method consistently outperforms other comparative methods on both datasets, achieving average mIoU improvements of 1.3% and 2.4%, respectively. Notably, the proposed method demonstrates even greater gains on smaller sampling splits, such as 1% and 10%, which are particularly beneficial for the semi-supervised setting. Specifically, with a 1% ratio, our method shows mIoU improvements of up to 1.9% and 5.3% on the two datasets, respectively, compared to the recently proposed SOTA method, DDSemi [21]. These results also validate that our method is less sensitive to reduced labeled data sampling than other methods. Furthermore, we present qualitative results in Fig. 5. As indicated by the purple and red dashed circles, our method achieves better segmentation of boundary regions and small objects.
@@ 4.4 Ablation Study
Figure 6: (a) Illustration of performance differences whether using point erasure strategy on the LaserMix [15]. (b) Variation in the proportion of points removed by the point erasure strategy throughout the training process. (c) Comparison of different mixing-based techniques. (d) Results of mIoU against the number of patch split $n$ under the 10% SemanticKITTI split.
@@ 5 Conclusion
This paper explores scene affinity for semi-supervised LiDAR semantic segmentation in driving scenes, named AIScene. Specifically, AIScene addresses intra-scene inconsistency through a point erasure strategy, which enhances the learning of pseudo-labeled points. Moreover, it employs a patch-based data augmentation method to mix multiple scenes at both scene and instance levels, guided by inter-scene correlation. This method enriches the semantic diversity of labeled scenes and complements pseudo-labeled scenes. Extensive experiments show that AIScene achieves state-of-the-art results on popular benchmarks, with remarkable improvements in the most challenging settings. Additionally, the novel erasure strategy and augmentation method can function independently to facilitate semi-supervised LiDAR semantic segmentation.

### v3

@@ Abstract
This paper explores scene affinity (AIScene), namely intra-scene consistency and inter-scene correlation, for semi-supervised LiDAR semantic segmentation in driving scenes. Adopting teacher-student training, AIScene employs a teacher network to generate pseudo-labeled scenes from unlabeled data, which then supervise the student network’s learning. Unlike most methods that include all points in pseudo-labeled scenes for forward propagation but only pseudo-labeled points for backpropagation, AIScene removes points without pseudo-labels, ensuring consistency in both forward and backward propagation within the scene. This simple point erasure strategy effectively prevents unsupervised, semantically ambiguous points (excluded in backpropagation) from affecting the learning of pseudo-labeled points. Moreover, AIScene incorporates patch-based data augmentation, mixing multiple scenes at both scene and instance levels. Compared to existing augmentation techniques that typically perform scene-level mixing between two scenes, our method enhances the semantic diversity of labeled (or pseudo-labeled) scenes, thereby improving the semi-supervised performance of segmentation models. Experiments show that AIScene outperforms previous methods on two popular benchmarks across four settings, achieving notable improvements of 1.9% and 2.1% in the most challenging 1% labeled data. The code will be released at https://github.com/azhuantou/AIScene.
footnotetext: Equal contribution.  Corresponding authors.
@@ 1 Introduction
AIScene has been extensively evaluated through experiments and demonstrated superiority on two popular benchmarks for semi-supervised LiDAR semantic segmentation. It exceeds state-of-the-art methods by 1.3% on the SemanticKITTI [2] and 1.2% on the nuScenes [10] in terms of average performance across four settings (1%, 10%, 20%, and 50% labeled data). Notably, in the most challenging setting, with only 1% labeled data, AIScene achieves remarkable improvements of 1.9% and 2.1% on the two datasets. Additionally, our ablation studies confirm AIScene’s potential to reduce the requirements for labeled data across different LiDAR semantic segmentation networks (see Tab. 1 and  Tab. 2), with each component of AIScene able to function independently to facilitate semi-supervised LiDAR semantic segmentation (see Fig. 6).
@@ 2 Related Work
3D Point Cloud Data Augmentation: Point cloud data augmentation has been explored across various 3D vision tasks [56, 34, 44, 9, 43, 25]. Inspired by the success of Mixup [51] in image classification, several methods have investigated techniques for mixing selected point cloud samples. To preserve structural information, RSMix [20] introduces a neighboring function, and PointCutMix [52] establishes correspondences for sample fusion. Notably, these methods focus on local object-level augmentation. In response, several works have explored mixing strategies for point cloud scene augmentation. PloarMix [47] swaps points within the fan-shaped areas between two scenes, while CoSMix [32] aims to mitigate domain shift by mixing semantic point groups. Unlike these methods, our method is capable of modeling inter-scene information across multiple scenes to create more diversified data, while also extracting instance-level semantic information for individual classes to further enrich scenes.
@@ 3.3 Patch-based Data Augmentation
Following the same steps used for labeled scenes, we further create a pseudo patch pool $\mathcal{B}_{\text{Pat}}^{u}$ and a pseudo instance pool $\mathcal{B}_{\text{Ins}}^{u}$ from the erased scene $\hat{x}_{i}^{u}$. Unlike $\mathcal{B}_{\text{Pat}}^{l}$ and $\mathcal{B}_{\text{Ins}}^{l}$ covering the whole training set $D_{l}$, the pseudo pools $\mathcal{B}_{\text{Pat}}^{u}$ and $\mathcal{B}_{\text{Ins}}^{u}$ only contain scene and instance information from the current batch and are reset in each iteration. This prevents low-quality pseudo labels from persisting in pools and misleading model training. Additionally, if the number of points within a patch or instance falls below a certain threshold $\tau_{\text{min}}$, it will not be saved into the pool. In practice, we set $\tau_{\text{min}}$ to 5 [49]. If a patch or instance has fewer than 5 points, its spatial structure exhibits high uncertainty, making it semantically ambiguous. Thus, it is excluded from the pools to prevent noise during data augmentation.
@@
Ours (AIScene) - 61.2 66.3 67.4 67.9 65.7 60.2 72.3 75.0 76.4 70.9
Improvement $\uparrow$ - +1.9 +1.2 +1.1 +0.9 +1.3 +2.1 +2.1 +1.0 -0.1 +1.2
Table 1: Experimental results on the SemanticKITTI and nuScenes dataset compared with different SSL methods. The results are reported with mIoU scores (%). The bold and underline scores respectively denote the best and second results.
@@ 4.2 Implementation Details
We conduct experiments with 4$\times$ NVIDIA V100 GPUs with 32GB RAM, based on MMDetection3D [8]. For the threshold $\tau_{s}$ for pseudo-label generation and point erasure, we experimentally find that 0.9 yields relatively satisfactory results. We use the MinkowskiNet [7] and Cylinder3D [57] as the backbone network due to its widely adopted in previous methods [15, 28, 42]. To facilitate patch-based data augmentation, we restrict our focus to point clouds within the X and Y dimensions in the range of $[-50,50]m$ for both the SemanticKITTI and nuScenes datasets. Similar to mainstream work, we also adopt the consistency loss proposed in MeanTeacher [39] to align the predictions of the teacher and student networks. Additionally, the weights of the consistency loss are assigned as $1\times 10^{3}$ and $2\times 10^{3}$ for the SemanticKITTI and nuScenes datasets, respectively. The weights for $\lambda_{u}$ and $\lambda_{l}$ are set as 1 for both datasets. Our method excludes all forms of model ensemble and test-time augmentation during the evaluation.
@@ 4.3 Main Results
Tab. 1 summarizes the quantitative results on SemanticKITTI and nuScenes datasets. Our method consistently outperforms other comparative methods on both datasets, achieving average mIoU improvements of 1.3% and 1.2%, respectively. Notably, the proposed method demonstrates even greater gains on smaller sampling splits, such as 1% and 10%, which are particularly beneficial for the semi-supervised setting. Specifically, with a 1% ratio, our method shows mIoU improvements of up to 1.9% and 2.1% on the two datasets, respectively, compared to the recently proposed SOTA method, DDSemi [21]. These results also validate that our method is less sensitive to reduced labeled data sampling than other methods. Furthermore, we present qualitative results in Fig. 5. As indicated by the purple and red dashed circles, our method achieves better segmentation of boundary regions and small objects.
@@ 4.4 Ablation Study
Figure 6: (a) Illustration of performance differences whether using point erasure strategy on the LaserMix [15]. (b) Variation in the proportion of points removed by the point erasure strategy throughout the training process. (c) Comparison of different mixing-based techniques on SemanticKITTI dataset. (d) Results of mIoU against the number of patch split $n$ under the 10% SemanticKITTI split.
@@ 5 Conclusion
This paper explores scene affinity for semi-supervised LiDAR semantic segmentation in driving scenes, named AIScene. Specifically, AIScene addresses intra-scene inconsistency through a point erasure strategy, which enhances the learning of pseudo-labeled points. Moreover, it employs a patch-based data augmentation method to mix multiple scenes at both scene and instance levels, guided by inter-scene correlation. This method enriches the semantic diversity of labeled scenes and complements pseudo-labeled scenes. Extensive experiments show that AIScene achieves state-of-the-art results on popular benchmarks. Additionally, the novel erasure strategy and augmentation method can function independently to facilitate semi-supervised LiDAR semantic segmentation.
Acknowledgments. This work is supported by the National Natural Science Foundation of China (No.62325111, U22B2011, 62271354).